# Improved Anonymous Multi Agent Path Finding Algorithm

### 17 Dec 2023 v1

@@ Introduction
We note that a similar but a more specific concept (in time-spaces), was used in  (Phillips and Likhachev 2011)(Gonzalez, Dornbush, and Likhachev 2012) to implicitly compress and expand the states generated by the wait actions of agents  [^1] The mentioned algorithms were originally tailored to graph-based pathfinding in the presence of dynamic obstacles where each vertex of the graph had to be annotated with the (safe) time intervals and the search utilized implicit move-then-wait actions. Our suggested algorithm on the other hand handles explicit graphs, and the notion of connected sequence in general is not obliged to relate to time dimension, as will be shown later in the paper.. Let us assume that while traversing the graph naively by single nodes, we are to expand a state $(v,h)$ which is located in a connected-sequence $(v,[h_{min},h_{max}])$. We can note that the nodes with the higher heights inside the connected-sequence (i.e. the nodes $(v,x):x\in[h+1,h_{max}]$) are all achievable from $(v,h)$ via the wait and restriction edges. Hence, the idea is to directly (and implicitly) generate all of these states ($(v,x):x\in[h+1,h_{max}]$), form an implicit bulk consisting of these states (including the originally picked-up node $(v,h)$), and expand them all at once. We will refer to the described mechanism of generating the sequential successors of states that uses only wait and restriction edges as to the straightforward generation. Note that it is enough to store the vertex and the height bounds of the bulk to define it. So technically when forming a bulk (for future expansion) we only need to find the last straightforwardly achievable state (i.e. the highest state in the connected-sequence).
@@
When the bulk is ready, it is expanded in the following fashion. First, let us assume the naive expansion of a bulk when for every node that resides in it we generate all of its immediate successors. Now observe that as a result of such procedure we are likely to have numerous successors that are characterized by the same graph vertex and different heights. Moreover many of these successors may belong to the same connected sequence. Thus, instead of explicit generation of them we generate only the ones with lower heights in their sequences. As the other search states from these sequences will be straightforwardly generated later on (i.e. when the search state with the lowest copy will be picked for expansion and its bulk will be formed as described above).
@@
In other words, to expand a bulk of states $(v,[h_{l},h_{u}])$, we do the following. We iterate over all connected-sequences in neighbor vertices of $v$, in which we can achieve at least one node (from a node $(v,x):x\in[h_{l},h_{u}]$). Then, we only generate the accessible node with the minimum height in each of these connected-sequences. The periodic structure of the reduced networks allows us to find the node with minimal height fast (as we will show later). As the number of connected-sequences is much less than the number of individual nodes, this leads to high reduction in the number of generated states. We call the algorithm that utilizes the described concepts Bulk-Search (BS). Next, we describe the details of the implementation.
@@
Algorithm 1 shows the pseudo-code of a graph traversal algorithm with our modifications i.e. bulk search. First we order the states inside the search set (OPEN) by their height i.e. we always choose the state with the minimum height for the expansion. This will help us in reducing the number of expansions as we will see later. Second change is that whenever we need to check whether a chosen state was expanded before, we additionally check if the state can be straightforwardly generated from the previously opened states. In this case, we do not need to expand it, as we assume that it was implicitly expanded before, and its successors were already generated and inserted into the search set. This can be done by checking if any state in the same connected-sequence and lower height was expanded before (i.e. stored in CLOSED set)(lines 11-14). This check should also be done when inserting new states into OPEN set (lines 20-23).
@@
The third change is how we generate the successors of all states (the taken one from OPEN along with its straightforwardly generated states) fast. The pseudo-code is presented in Algorithm 2. Firstly (lines 1-3), if the node is the source node then we have only one node as input (i.e. no implicit states) and thus we need to only generate neighboring successors (i.e. connected nodes in the network) and return them. Otherwise, if the input node $n(v,h)$ is an inner node in the connected-sequence $(v,[h_{min},h_{max}])$, we should generate the successors of all states $n(v,x):x\in[h,h_{max}]$. This can be done as follows. If the state $(v,h)$ is located in the beginning of the connected-sequence (i.e. $h=h_{min}$), this state may have reversed edges, so we always generate all neighboring nodes of this state in the network. The same thing is applied on $(v,h_{max})$ where we also generate all its neighboring nodes in the network. Other nodes (i.e. nodes $(v,x):x\in[max(h_{min}+1,h),h_{max}-1]$) have only move edges to connect to nodes in other connected-sequences, so we can do the following (lines 10-17) to generate their successors. We iterate over all connected-sequences $cs$ in neighbor vertices of $v$. We then check whether we can achieve at least one node in $cs$ (from nodes $(v,x):x\in[max(h_{min}+1,h),h_{max}-1]$). As shown in lines 11-13, this can be done by checking whether there is at least one move edge which starts from an outer copy in $[max(h_{min}+1,h),h_{max}-1]$ and ends at inner copy in $cs$. If so, we generate the node with the minimum accessible height in $cs$ and added to successors set (lines 14-15).
@@
As a result, any achievable node from the source node is either explicitly or implicitly (from a node with lower height in the same connected-sequence) expanded. Therefore, we can immediately state the following theorem.
@@
Figure 2: Example showing the connected-sequences on the network. Yellow crossbars denote to the connected-sequences on each vertex. Initially, we have the connected-sequences as shown in the upper figure. After a path (green one) is found and its edges are reversed, the connected-sequences are divided as shown in the lower figure.
@@
The search states inside the search set are sorted according to the height. This helps in reducing the number of expansions as lower height states implicitly expands the higher height states in the same connected-sequence but the opposite is not applicable. Theoretically multiple single states in one connected-sequence may be expanded individually e.g. if they opened in descending-height way. This is possible even if we order the states according to heights as we have reversed edges which generates states with lower height. However, in practice only few states are expanded in each connected-sequence and thus the algorithm mainly depends on the number of connected-sequences. Fortunately, the number of connected-sequences is much smaller than the size of the network. Initially, we have a number of connected-sequences equals to the number of original graph vertices $|V|$. After each path is found, $T$ (the length of the path) new connected-sequences appear. As a result, in the whole search for all $k$ agents, the number of appeared connected-sequences equal $\sum_{i=1}^{i=k}{|V|+T(i-1)}=k|V|+Tk(k-1)/2$. Therefore, we have a theoretical reduction in the number of nodes (comparing with $k|V|T$, the number of nodes without compressing in connected-sequences) equals $min(|V|/k,T/2)$. This reduction is significantly high that allows us to obtain the fastest full success (to our knowledge) in optimally solving all public MAPF benchmarks for a MAPF-family problem, as will be shown in next section.
@@
Table 1: Succes rates of flow-BS and flow solvers with a timeout of 30 seconds.
@@
Figure 4: The number of expanded nodes with different number of agents on different maps.
@@
Figure 5: The figure shows for specific map and specific number of agents, how the number of expansions changes for increasing values of $T$ (the maximum height (copy)). The values of $T$ are chosen relative to the value of the real makespan $T_{min}$.
@@
We implemented the improved AMAPF solver in C++ [^2] https://github.com/PathPlanning/AMAPF-with-MF-and-Bulk-Search.git and compare it with the state-of-the-art AMAPF solver that does not use the introduced bulk search to solve MF but rather utilizes the standard Ford-Fulkerson as suggested in (Yu and LaValle 2013b). The code of the competitor was taken from public repo  [^3] https://github.com/Kei18/tswap that accompanied the paper (Okumura and Défago 2022). We kept all the optimization techniques designed by the code authors. We will denote these two solvers as flow-BS (ours) and flow (state of the art). The experiments were conducted on a PC with Intel Core i7-10700F CPU @ 2.90GHz $\times$ 16 and 32Gb of RAM.
@@
The MAPF maps and instances were taken from the publicly available MAPF benchmark (Stern et al. 2019b). We use all 33 maps available in this benchmark and all 25 random scenarios. Each scenario on each map (except some small ones), contains 1000 pairs of start-goal positions. To test a solver on a scenario, we run it with the first 1, 2, 4, 8, 16, 32, 64, 128, 256, 512, 1000 pairs, sequentially. Whenever a solver fails to solve a problem under a time limit of 30 seconds, we terminate testing on this scenario and move to the next one.
@@
In the first experiment we used a precise estimator of $T$ suggested by (Okumura and Défago 2022) (which solves bottleneck assignment problem (Gross 1959)) to estimate the lower bound of the makespan. As this estimator takes non-negligible time (up to 10 seconds) and both algorithms use it, we did not account its time from the 30s time budget.
@@
For each map, we collected the total number of success instances over all scenarios. The table 1 summarizes the results. As can be shown, our solver was able to solve all instances in all maps under 30s. However, this was not the case for flow solver. The latter searches the network node-by-node and, therefore, was able to solve accomplish test only on the the small maps. When maps are large (like city maps that are $256\times 256$) or the makespan is high (like in the maze maps), it often was not able to produce solution under the imposed limit. Fig. 3, shows the total success rate for all instances in all maps every one second. In fact, flow-BS was able to solve the hardest instance (orz900d map, scenario 20 with 256 agents) in 17s. On the other hand, flow showed very small increase of SR after solving the easy instances ( 75% of instances) around the 8th second.
@@
In the second experiment, we investigated how the number of agents affects the performance. For this purpose, we selected three maps of different topology and size and plotted the average number (over all scenarios) of expanded nodes (machine-independent metric) while searching, against the number of agents. The results (plotted in Fig. 4) show that our algorithm expands much less nodes than standard algorithm (as expected). This again proves the dependence of our algorithm on the number of connected-sequences in the network but not the network size. Moreover, the number of connected-sequences increases linearly with the number of agents (it approaches to quadratically when the number of agents approaches the size of the network) as was deduced theoretically earlier in this paper. For standard Ford-Fulkerson algorithm, the search also increases linearly with the number of agents, but with the size of the whole network. It worth to note here, that the relation between the runtime and number of expansions is not fixed but also increases. This is because the single memory-access/read/write operation takes more time when the amount of the stored data increases.
@@
So far, we assumed that we have a good estimator which can estimate an optimal or near-optimal lower bound ($T$) of the makespan. However this may be not the case, therefore the search may be repeated several times until finding the optimal solution. Therefore, we also conducted experiments to show the practical performance of both solvers for different $T$ when we fix the map and the number of agents. The tests were designed as follows. For each one of three fixed maps (the maps warehouse-20-40-10-2-2, maze-128-128-10 and room-64-64-16 were chosen), and for a number of agents $\in\{256,1000\}$, we run the solvers on networks with maximum height (copy) equal one of the values $\{T_{min}/4,T_{min}/2,T_{min},T_{min}*2\}$, where $T_{min}$ is the optimal makespan. Again, we recorded the average number of expanded nodes over all scenarios (see Fig. 5). Obviously, the number of nodes expanded by flow significantly increases with the value of $T$, while flow-BS does not demonstrate such growth. That means that our solver is especially useful when one can not estimate the optimal makespan accurately before actually solving an AMAPF problem instance.
@@ Conclusion
In this paper we have revisited the reduction-based approach to optimally solving Anonymous MAPF problem, when the latter is reduced to a search problem on an auxiliary graph of a special structure. We have suggested an improved AMAPF solver that is based on a specific search algorithm, tailored to find paths on the auxiliary graphs exploiting their specific topology. We have showed that our improved AMAPF solver significantly outperforms state of the art on a large variety of setups, due to its better scalability to the size of the input graph.

### 19 Dec 2023 v2

@@ Introduction
We note that a similar but more specific concept was used in  (Phillips and Likhachev 2011) and (Gonzalez, Dornbush, and Likhachev 2012) to implicitly compress and expand the states generated by the wait actions of agents  [^2] The mentioned algorithms were originally tailored to graph-based pathfinding in the presence of dynamic obstacles where each vertex of the graph had to be annotated with the (safe) time intervals and the search utilized implicit move-then-wait actions. Our algorithm on the other hand handles explicit graphs, and the notion of connected-sequence in general is not obligated to relate to the time dimension, as will be shown later in the paper.. Let us assume that while traversing the graph naively by single nodes, we are to expand a state $(v,h)$ located in a connected-sequence $(v,[h_{min},h_{max}])$. We can note that the nodes with the higher heights inside the connected-sequence (i.e. the nodes $(v,x):x\in[h+1,h_{max}]$) are all achievable from $(v,h)$ via the wait and restriction edges. Hence, the idea is to directly (and implicitly) generate all of these states ($(v,x):x\in[h+1,h_{max}]$), form an implicit bulk consisting of these states (including the originally picked-up node $(v,h)$), and expand them all at once (instead of expanding only $(v,h)$). We will refer to the described mechanism of generating the sequential successors of states that uses only wait and restriction edges as straightforward generation. Note that it is enough to store the vertex and the height bounds of the bulk to define it. So, technically, when forming a bulk (for future expansion), we only need to find the last straightforwardly achievable state (i.e. the highest state in the connected-sequence).
@@
When the bulk is ready, it is expanded in the following fashion. First, let us assume the naive expansion of a bulk when, for every node that resides in it, we generate all of its immediate successors. Now observe that, as a result of such procedure, we are likely to have numerous successors that are characterized by the same graph vertex and different heights. Moreover, many of these successors may belong to the same connected-sequence. Thus, instead of explicitly generating them, we generate only the ones with lower heights in their sequences. The other search states from these sequences will be straightforwardly generated later on (i.e., when the search state with the lowest copy is picked for expansion and its bulk is formed as described above).
@@
In other words, to expand a bulk of states $(v,[h_{l},h_{u}])$, we do the following. We iterate over all connected-sequences in neighbor vertices of $v$, in which we can achieve at least one node (from a node $(v,x):x\in[h_{l},h_{u}]$). Then, we only generate the accessible node with the minimum height in each of these connected-sequences. The periodic structure of the reduced networks allows us to quickly find the node with the minimal height (as we will show later). As the number of connected-sequences is much less than that of individual nodes, it leads to a high reduction in the number of generated states. We call the algorithm that utilizes the described concepts Bulk Search (BS). Next, we describe the details of the implementation.
@@
Algorithm 1 shows the pseudo-code of a graph traversal algorithm with our modifications. First, we order the states inside the search set (OPEN) by their height, i.e., we always choose the state with the minimum height for the expansion. This will help us reduce the number of expansions as will be shown later. The second change is that whenever we need to check whether a chosen state was expanded before, we additionally check if the state can be straightforwardly generated from the previously opened states. In this case, we do not need to expand it, as we assume that it was implicitly expanded before and its successors were already generated and inserted into the search set. This can be done by checking if any state in the same connected-sequence and lower height was expanded before (i.e. stored in the CLOSED set) (lines 11-14). This check should also be done when inserting new states into the OPEN set (lines 20-23).
@@
The third change is how we generate the successors of all states (the taken one from OPEN along with its straightforwardly generated states) fast. The pseudo-code is presented in Algorithm 2. Firstly, (lines 1-3), if the node is the source node $s$, we have only one node as input (i.e. no implicit states), and, thus, we need to only generate neighboring successors (i.e. connected nodes in the network) and return them. Otherwise, let the input node $n(v,h)$ (which is not $s$ or $g$) located in the connected-sequence $(v,[h_{min},h_{max}])$, then we should generate the successors of all states $n(v,x):x\in[h,h_{max}]$. This can be done as follows. If the state $(v,h)$ is located at the beginning of the connected-sequence (i.e. $h=h_{min}$), this state may have reversed edges, so we always generate all neighboring nodes of this state in the network. The same thing is applied on $(v,h_{max})$ where we also generate all its neighboring nodes in the network. Other nodes (i.e. nodes $(v,x):x\in[max(h_{min}+1,h),h_{max}-1]$) have only move edges to connect to nodes in other connected-sequences, so we can do the following (lines 10-17) to generate their successors. We iterate over all connected-sequences $cs$ in neighbor vertices of $v$. We then check whether we can achieve at least one node in $cs$ (from nodes $(v,x):x\in[max(h_{min}+1,h),h_{max}-1]$). As shown in lines 11-13, this can be done by checking whether there is at least one move edge which starts from an outer copy in $[max(h_{min}+1,h),h_{max}-1]$ and ends at inner copy in $cs$. If so, we generate the node with the minimum accessible height in $cs$ and add it to the successors set (lines 14-15).
@@
As a result, any achievable node beginning from the source node is either explicitly or implicitly (from a node with lower height in the same connected-sequence) expanded. Therefore, we can immediately state the following theorem.
@@
Figure 2: Example showing connected-sequences on the network. Yellow crossbars denote the connected-sequences on each vertex. Initially, we have the connected-sequences as shown in the upper figure. After a path (green one) is found and its edges are reversed, the connected-sequences are divided as shown in the lower figure.
@@
The search states inside the search set are sorted according to the height. This helps to reduce the number of expansions as the lower-height states implicitly expand the higher-height states in the same connected-sequence, but the opposite is not applicable. In theory, multiple single states in one connected-sequence may be expanded individually, e.g., if they have been opened in the descending-height way. This is possible even if we order the states according to the heights as we have reversed edges which generates states with a lower height. However, in practice, only few states are expanded in each connected-sequence, and, thus, the algorithm’s performance mainly depends on the number of connected-sequences. Fortunately, the number of connected-sequences is much smaller than the size of the network. Initially, we have a number of connected-sequences equals $V$, the number of original graph vertices. After each path is found and reversed, $T$ new reversed edges, and therefore, $T$ new connected-sequences appear. As a result, in the whole search for all $k$ agents, the number of the appearing connected-sequences equals $\sum_{i=1}^{i=k}{|V|+T(i-1)}=k|V|+Tk(k-1)/2$. Therefore, we have a theoretical reduction in the number of nodes (comparing with $k|V|T$, the number of nodes without compressing in connected-sequences) equals $min(|V|/k,T/2)$. This reduction is significantly high, which allows us to obtain the fastest full success (to our knowledge) in optimally solving all public MAPF benchmarks for a MAPF-family problem, as will be shown in the next section.
@@
Table 1: Success rates of flow-BS and flow solvers with a timeout of 30 seconds.
@@
Figure 4: The number of expanded nodes with a different number of agents on different maps.
@@
Figure 5: The figure shows a specific map and a specific number of agents, how the number of expansions changes for increasing values of $T$ (the maximum height (copy)). The values of $T$ are chosen relative to the value of the real makespan $T_{min}$.
@@
We have implemented the improved AMAPF solver in C++ [^3] https://github.com/PathPlanning/AMAPF-MF-BS and compared it with the state-of-the-art AMAPF solver that does not use the introduced Bulk Search to solve MF but rather utilizes the standard Ford-Fulkerson as suggested in (Yu and LaValle 2013b). The code of the competitor was taken from public repository  [^4] https://github.com/Kei18/tswap that accompanied the paper (Okumura and Défago 2022). We kept all the optimization techniques designed by the code authors. We will denote these two solvers as flow-BS (ours) and flow (state of the art). The experiments were conducted on a PC with Intel Core i7-10700F CPU @ 2.90GHz $\times$ 16 and 32Gb of RAM.
@@
The MAPF maps and instances were taken from the publicly available MAPF benchmark (Stern et al. 2019b). We use all 33 maps available in this benchmark and all of the 25 random scenarios. Each scenario on each map (except some small ones) contains 1,000 pairs of start-goal positions. To test a solver on a scenario, we run it with the first 1, 2, 4, 8, 16, 32, 64, 128, 256, 512, and 1,000 pairs of start-goals sequentially. Whenever the solver fails to solve a problem under a time limit of 30 seconds, we terminate testing on this scenario and move to the next one.
@@
In the first experiment, we have used a precise estimator of $T$ suggested by (Okumura and Défago 2022) (which solves bottleneck assignment problem (Gross 1959)) to estimate the lower bound of the makespan. As this estimator takes non-negligible time (up to 10 seconds) and both algorithms use it, we have not accounted for its runtime.
@@
For each map, we have collected the total number of success instances over all scenarios. Table 1 summarizes the results. As can be seen, our solver is able to solve all instances in all maps under 30s. However, this is not the case for flow solver. The latter searches the network node-by-node and, therefore, is able to solve the test instances only on the small maps. When the maps are large (like city maps that are $256\times 256$) or the makespan is high (like in the maze maps), it is often unable to produce a solution under the imposed time limit. Fig. 3 shows the success rate for all instances in all maps second by second. In fact, flow-BS is able to solve the hardest instance (orz900d map, scenario 20 with 256 agents) in 17s. On the other hand, flow has shown a very slight increase of SR after solving the easy instances ( 75% of instances) around the 8th second.
@@
In the second experiment, we have investigated how the number of agents affects the performance. For this purpose, we selected three maps of different topology and size and plotted the average number of the expanded nodes against the number of agents. The results (plotted in Fig. 4) show that our algorithm expands much fewer nodes than the standard algorithm (as expected). This again proves the dependence of our algorithm on the number of connected-sequences in the network but not the network size. Moreover, the number of connected-sequences increases linearly with the number of agents (it approaches to quadratically only when the number of agents approaches the size of the network) as has been deduced theoretically earlier in this paper. For the standard Ford-Fulkerson algorithm, the search also increases linearly with the number of agents, but with a factor of the size of the whole network. It is worth noting that the relation between the runtime and the number of expansions is not fixed but increases for $flow$, too. This is because a single memory-access/read/write operation takes longer when the amount of the stored data increases.
@@
So far, we have assumed that we have an estimator which can identify an exact bound ($T$) of the makespan. However, this may be not the case, so the search may be repeated several times until finding the optimal solution. Therefore, we have also conducted experiments to show the practical performance of both solvers for different $T$ when we fix the map and the number of agents. The tests are designed as follows. For each one of the three fixed maps, warehouse-20-40-10-2-2, maze-128-128-10, room-64-64-16, and for a number of agents $\in\{256,1000\}$, we run the solvers on networks with a maximum height equals one of the values $\{T_{min}/4,T_{min}/2,T_{min},T_{min}*2\}$, where $T_{min}$ is the optimal makespan. Again, we have recorded the average number of expanded nodes over all scenarios (see Fig. 5). Obviously, the number of nodes expanded by flow significantly increases with the value of $T$, while flow-BS does not demonstrate such growth. It means that our solver is especially useful when one can not estimate the optimal makespan accurately before actually solving an AMAPF problem instance.
@@ Conclusion
In this paper, we have revisited the reduction-based approach to optimally solving the Anonymous MAPF problem, when the latter is reduced to a search problem on an auxiliary graph of a special structure. We have suggested an improved AMAPF solver that is based on a specific search algorithm tailored to find paths on the auxiliary graphs exploiting their specific topology. We have shown that our improved AMAPF solver significantly outperforms the state-of-the-art one on a large variety of setups, leveraging its better scalability to the size of the input graph. Next, we plan to extend the proposed search technique to support costs (e.g. Min-Cost-Maximum-Flow) to solve other MAPF problems.